OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

OpenStudy (anonymous):



OpenStudy (anonymous):

\[| x - 5 |\ge 3\] \[x \le 2\] \[x \ge 8\] \[X \epsilon (- \infty , 2]\upsilon [8 , \infty)\]